The perimeter of a rectangle if 68 feet. If the length is 8 feet more than the width, find both the length and width. (Formula: P = 2L + 2W)
Answer by algebrahouse.com(1659)

x = width
x + 8 = length {the length is 8 more than the width}

Perimeter of a rectangle is 2(width) + 2(length)

2x + 2(x + 8) = 68 {substituted expressions into perimeter formula}
2x + 2x + 16 = 68 {used distributive property}
4x + 16 = 68 {combined like terms}
4x = 52 {subtracted 16 from each side}
x = 13 ft width {divided each sideb y 4}
x + 8 = 21 ft length {substituted 13, in for x, into x + 8}

width is 13 ft
length is 21 ft

For more help from me, visit: www.algebrahouse.com